# Six magnetization plateau phases in a kagome antiferromagnet with $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$-type distortion

### v1

@@ Abstract
A recently discovered kagome antiferromagnet $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$ has attracted significant interest due to its unique kagome lattice structure and magnetic properties. Two separate studies have proposed distinct sets of exchange interaction parameters for this compound. In this study, we investigate the ground state of the spin-1/2 Heisenberg kagome model with three types of nearest-neighbor exchange interactions under a magnetic field by exact diagonalization using the Lanczos method. We reveal that clear magnetization plateaus at $M/M_{\rm sat}$=1/3, 5/9, and 7/9 are present under both parameter sets, which are identified as magnon crystal states based on their spin structures. Our findings suggest that these plateaus could potentially be experimentally confirmed with magnetization measurements on $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$ under a magnetic field of approximately 300 T. Due to the uncertainty in the exchange interactions of $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$, we perform additional calculations by varying these interactions. Consequently, we discover additional plateau phases at $M/M_{\rm sat}$=1/3, 5/9, and 7/9, each distinctly different from the magnon crystal states.
@@ I Introduction
Quantum kagome antiferromagnets are renowned for hosting novel ground states due to the influence of frustration, garnering significant attention in the field of condensed matter physics [1, 2, 3]. For the case of $S=1/2$, under zero magnetic field, various theoretical calculations predict the existence of quantum spin liquids with finite gaps between singlet and triplet states [4, 5, 6], gapless quantum spin liquids [7, 8, 9, 10, 11] and valence bond crystal (VBC) states [12, 13, 14]. Moreover, under a magnetic field, theoretical investigations employing techniques such as diagonalization and the density matrix renormalization group method have revealed the emergence of magnetization plateaus. These plateaus occur at $M/M_{\rm sat}$=1/9, 1/3, 5/9, and 7/9, where $M$ represents magnetization and $M_{\rm sat}$ denotes saturation magnetization [15, 16, 17, 18, 19, 20, 21, 22]. Among them, the 7/9 plateau is rigorously established to exist, corresponding to a state referred to as the magnon crystal state (MCS) or VBC state [23], where magnons are periodically localized on the lattice. Similarly, it is anticipated that the 1/3 and 5/9 plateaus also correspond to the MCS [16, 17, 24, 25]. The state of the 1/9 plateau remains theoretically contentious, predicted to be as either a $Z_{3}$ spin liquid [17] or a VBC [18, 21] state, with detailed magnetic structure yet to be elucidated. Very recently and interestingly, the 1/9 and 1/3 plateaus have also been experimentally observed in the kagome compound $\rm YCu_{3}(OH)_{6+\it x}Br_{3-\it x}$ ($x\approx 0.5$[26, 27, 28]. Therefore, theoretical and experimental studies on magnetization plateaus in the kagome lattice are currently attracting significant attention. These theoretical studies have primarily focused on the ideal kagome lattice where all exchange interactions are equivalent. However, in many kagome antiferromagnetic compounds, distorted kagome lattices are formed, resulting in non-equivalent exchange interactions [29]. Nevertheless, theoretical investigations of distorted kagome lattices are scarce [30, 31]. Therefore, it can be expected that new quantum states, which do not exist in undistorted kagome lattices, may emerge.
@@
$\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$ is a recently synthesized compound where copper ions ($\rm Cu^{2+}$) form a distorted kagome lattice [32, 33, 34, 35, 36, 37]. There are three types of nearest-neighbor exchange interactions. A previous study has theoretically revealed the existence of a ground state with $Q=(1/3,1/3)$, coplanar long-range order in the classical spin system under zero magnetic field, which has also been experimentally observed by means of neutron scattering experiments. The candidate values for the exchange interactions have been proposed in two previous studies [35, 36]. In this paper, we refer to them as Type 1 [35] and Type 2 [36]. However, the quantum magnetic properties of the quantum spin system under the magnetic field have not been sufficiently investigated, and thus remain unclear. Therefore, it is important to investigate the ground state in the quantum spin model for $\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$.
@@
In this study, we investigate the ground state of this model under the magnetic field by exact diagonalization using the Lanczos method. We reveal that clear 1/3, 5/9, and 7/9 plateaus are present for both the Type 1 and Type 2 parameter sets. These ground states are identified as the MCS based on their spin structures. Given the uncertainty in the exchange interactions of $\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$, we perform additional calculations by varying the exchange interactions within a certain range. Consequently, we find additional plateau phases at $M/M_{\rm{sat}}=1/3$, 5/9, and 7/9, each distinctly different from the MCS. We argue that clear magnetization plateaus can be observed through high-field magnetization measurements conducted on $\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$.
@@ II Model
We perform exact diagonalization using the Lanczos method for clusters with $N=27$, 36, and 45 under periodic boundary conditions (PBCs) as shown in Fig. 1, where $N$ represents the number of sites.
@@ III.1 Magnetization curve of $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$
Figure 2 depicts the magnetization curves at zero temperature of the Type 1 and Type 2 for ${\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}}$. Here, we set $g=2$. Both the Type 1 and Type 2 exhibit clear plateaus at $M/M_{\rm{sat}}=$ 1/3, 5/9, and 7/9. In the Type 1, the widths of these plateaus show almost no dependence on the number of sites $N$, indicating negligible finite-size effects. Thus, these plateaus persist even in the thermodynamic limit. In the Type 2, while the effects of size are negligible for the 5/9 and 7/9 plateaus, the width of the 1/3 plateau varies significantly at $N=27$ and $N=36$. However, the plateau width is broader for $N=36$ than for $N=27$, indicating its significant persistence even in the thermodynamic limit. The magnetic structures of all three magnetization plateaus are confirmed to be the MCS in Sec. III.2. Figure 2 shows that all of these magnetization plateaus can be observed through magnetization measurements on $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$. Notably,the 1/3 plateau is observable at approximately 150 T, and the possibility of observing all plateaus increases at a stronger field of approximately 300 T.
Figure 2: Magnetization curves of the kagome lattice with (a) the Type 1 and (b) Type 2 parameters. The blue dashed, red solid, and green solid lines represent the calculation results for $N=27$, 36, and 45, respectively. The black solid lines show the calculation results for the classical spin system at $N=5184$.
@@
Figure 4: Magnetization curves for the kagome lattice with the exchange interactions $J_{\rm r}$, $J_{\rm g}$, and $J_{\rm b}$, where $J_{\rm r}$ and $J_{\rm b}$ are fixed at 1, and $J_{\rm g}$ varies from 0 to 2. The blue dashed, red solid, and green solid lines represent the calculation results for $N=27$, 36, and 45, respectively.
In Fig. 2(a), the width of the 1/6 plateau at $N=36$ is comparable to that of the 1/3 plateau. Hence, the presence of the 1/6 plateau is also suggested. However, the 1/6 plateau is not observed in the $N=27$ system, and thus, there are insufficient results to conclude the existence of the plateau. The 1/6 plateau requires the number of sites to be a multiple of 12, and since the model has 9 sites per unit cell, $N$ must be a multiple of 36. Therefore, it is currently impossible to evaluate it through diagonalization.
Both the Type 1 and Type 2 show a magnetic jump immediately after $M/M_{\rm{sat}}=1/6$ at $N=36$. This jump occurs from $M/M_{\rm{sat}}$=3/18(1/6) to $M/M_{\rm{sat}}$=5/18, skipping $M/M_{\rm{sat}}$=4/18, suggesting the possibility of a first-order field-induced quantum phase transition. However, this is not observed at $N=27$. Thus, the jump observed at $N=36$ might be special due to finite-size effects. Further calculations with larger clusters are also necessary.
@@ III.2 Spin structures of the magnetization plateaus
respectively, while the thickness of the lines corresponds to their magnitude. The diameter of the magenta lattice points corresponds to the value of $\Braket{S_{i}^{z}}$ at that site. The spin correlations between bonds connected by the exchange interaction $J_{\rm r}$ are strongest, forming periodically arranged hexagons. The remaining spins align parallel to the magnetic field. This structure corresponds to the MCS. Therefore, the magnetic structures of these plateaus are the same as the anticipated magnetic structures of the plateaus arising in the uniform kagome lattice. The emergence of the 1/3, 5/9, and 7/9 plateaus can be understood by considering a model with only the $J_{\rm r}$ hexagon. In this model, magnetization plateaus occur at $M=0,1,2$. These plateaus, combined with the remaining spins aligned parallel to the magnetic field, give rise to the $M/M_{\rm sat}$=1/3, 5/9, and 7/9 plateaus on the kagome lattice.
@@ III.3 Calculations with other parameters
Two sets of exchange interaction parameters, the Type 1 and Type 2, listed in Table 1, have been proposed for $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$. This suggests that the parameters of this compound are not precisely determined at present. Therefore, analyzing the model based on other parameters is also essential for understanding its magnetism. Moreover, in the future, when new compounds are synthesized and their magnetic measurements are conducted, analyzing it based on alternative parameters would provide significant assistance. This is important from a theoretical perspective as well, given the possibility of discovering new magnetic phases. In the Type 1, $J_{\rm r}\approx J_{\rm b}\neq J_{\rm g}$ while in the Type 2, $J_{\rm r}=J_{\rm b}\neq J_{\rm g}$ as shown in Table 1. Therefore, in this section, we investigate the model with the restriction $J_{\rm r}=J_{\rm b}\neq J_{\rm g}$. We consider the normalized Hamiltonian (1) with $J_{\rm r}=J_{\rm b}=1$ and the magnetic field term $h=g\mu_{\rm B}H$.
@@
Figure 4 shows the magnetization curves for $0\leq J_{\rm g}\leq 2$. Interestingly, even at $J_{\rm g}=0$, the 1/3, 5/9, and 7/9 plateaus are observed, as evident from Fig. 4(a). Furthermore, for $J_{\rm g}<1$, these plateaus exhibit a broader width compared to that for $J_{\rm g}=1$. This suggests that these plateaus remain stable even when deviating significantly from the uniform kagome lattice ($J_{\rm g}=1$). Considering that the clear 5/9 and 7/9 magnetization plateaus have not yet been observed experimentally in the kagome lattice, the theoretical prediction of increased plateau widths provides sufficient motivation to conduct further experiments. For $J_{\rm g}>1$, the 5/9 and 7/9 plateau widths narrow as $J_{\rm g}$ increases. At $J_{\rm g}=1.5$, the plateau widths are significantly narrowed, as shown in Fig. 4(g), suggesting that these plateaus may not exist in the thermodynamic limit. In contrast, the width of the 1/3 plateau widens as $J_{\rm g}$ increases. At $J_{\rm g}=2$, as depicted in Fig. 4(h), it is evident that the width of the 1/3 plateau significantly widens. This is a natural consequence of the fact that when $J_{\rm g}$ becomes very large, it becomes equivalent to an isolated three-site model, resulting in the broadening of the 1/3 plateau.
@@
As explained above, the changing properties of the plateaus around $J_{\rm g}=1$ indicate the occurrence of a phase transition around $J_{\rm g}=1$. To confirm the presence or absence of phase transitions with respect to $J_{\rm g}$ at $M/M_{\rm sat}=1/3$, 5/9, and 7/9, we calculate the second derivative of the energy per site $e=E/N$ with respect to $J_{\rm g}$. Figure 5 shows the calculation results of $-\frac{d^{2}e}{dJ_{\rm g}^{2}}$ at $M/M_{\rm sat}=1/3$, 5/9, and 7/9. It is evident that all three results exhibit sharp peaks at $J_{\rm g}=1$. For $M/M_{\rm sat}=1/3$ and 5/9, based on the results in Figs. 5(a) and 5(b), it is not possible to precisely determine the dimension of the phase transition, but it is expected to be either second-order or first-order. Further calculations with larger system sizes are required to clarify the dimension of the phase transition. For $M/M_{\rm sat}=7/9$, as shown in Fig.5(c), it is apparent that $-\frac{d^{2}e}{dJ_{\rm g}^{2}}$ diverges at $J_{\rm g}=1$ sharply, indicating a first-order phase transition at $J_{\rm g}$.
@@
Figure 6: Results of finite-size scaling for the 5/9 plateau at $J_{\rm r}=J_{\rm b}=1$ and $J_{\rm g}=1.2$. The lower black line and upper black line represent the results of the linear fits for $h_{\rm l}$ and $h_{\rm u}$, respectively, which are defined as the lower and upper magnetic fields of the magnetization plateau. The red bars represent errors from the linear fitting.
@@
In the following we discuss the existence of plateaus in the thermodynamic limit. The plateau width for $M/M_{\rm sat}=1/3$ remains broad regardless of $J_{\rm g}$, suggesting the existence of plateaus in the thermodynamic limit. In contrast, for $M/M_{\rm sat}=5/9$ and $7/9$, particularly at $J_{\rm g}>1$, the narrow plateau widths, as depicted in Fig. 4, make it difficult to discern the presence of plateaus in the thermodynamic limit. Therefore, it is necessary to examine finite-size effects on the plateau width. Here, the lower magnetic field and upper magnetic field of the magnetization plateau are defined as $h_{\rm l}$ and $h_{\rm u}$, respectively. If $h_{\rm u}-h_{\rm l}$ remains positive in the thermodynamic limit, it implies the presence of plateaus. Figure 6 shows the result of finite-size scaling for the 5/9 plateau at $J_{\rm g}=1.2$. The black solid lines at the lower and upper sides correspond to $h_{\rm l}$ and $h_{\rm u}$, respectively. The red points represent $h_{\rm l}$ and $h_{\rm u}$ in the thermodynamic limit, and error bars from the linear fitting are also shown. Here, the errors for $h_{\rm l}$ and $h_{\rm u}$ are denoted as $\delta h_{\rm l}$ and $\delta h_{\rm u}$, respectively. If the relation
@@
is satisfied, it can be confidently concluded that the plateau exists. Table 2 and Table 3 present $W_{\rm min}$ for $M/M_{\rm sat}=5/9$ and 7/9, respectively, at various $J_{\rm g}$. For both $M/M_{\rm sat}=5/9$ and 7/9, $W_{\rm min}$ remains positive up to $J_{\rm g}=1.2$, indicating the presence of the plateaus for $J_{\rm g}\leq 1.2$.

### v3

# Six magnetization plateau phases in a spin-1/2 distorted kagome antiferromagnet: application to $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$

@@ Abstract
A recently discovered kagome antiferromagnet $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$ has attracted significant interest due to its unique kagome lattice structure and magnetic properties. The kagome lattice has three types of exchange interactions: one hexagonal coupling and two different triangular couplings. Previous studies have shown that its ground state is significantly different from that predicted for the undistorted kagome lattice, forming a coplanar spin state with a commensurate magnetic wave vector ${\mathbf{Q}}=(1/3,1/3)$. Two separate studies have proposed distinct sets of exchange interaction parameters for this compound. In this study, we investigate the ground state of the spin-1/2 Heisenberg kagome model with three types of nearest-neighbor exchange interactions under a magnetic field by exact diagonalization using the Lanczos method. We reveal that clear magnetization plateaus at $M/M_{\rm sat}$=1/3, 5/9, and 7/9 are present under both parameter sets, which are identified as magnon crystal states based on their spin structures. Our findings suggest that these plateaus could potentially be experimentally confirmed with magnetization measurements on $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$ under a magnetic field of approximately 300 T, achievable with state-of-the-art magnetic field generators. In order to get a deeper understanding of magnetism of $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$, we perform additional calculations by varying these interactions. Consequently, we discover additional plateau phases at $M/M_{\rm sat}$=1/3, 5/9, and 7/9, each distinctly different from the magnon crystal states.
@@ I Introduction
Quantum kagome antiferromagnets are renowned for hosting novel ground states due to the influence of frustration, garnering significant attention in the field of condensed matter physics [1, 2, 3]. For the case of $S=1/2$, under zero magnetic field, various theoretical calculations predict the existence of quantum spin liquids with finite gaps between singlet and triplet states [4, 5, 6], gapless quantum spin liquids [7, 8, 9, 10, 11] and valence bond crystal (VBC) states [12, 13, 14]. Moreover, under a magnetic field, theoretical investigations employing techniques such as exact diagonalization and the density matrix renormalization group methods have revealed the emergence of magnetization plateaus. These plateaus occur at $M/M_{\rm sat}$=1/9, 1/3, 5/9, and 7/9, where $M$ represents magnetization and $M_{\rm sat}$ denotes saturation magnetization [15, 16, 17, 18, 19, 20, 21, 22]. Among them, the 7/9 plateau is rigorously established to exist theoretically, corresponding to a state referred to as the magnon crystal state (MCS) or VBC state [23], where magnons are periodically localized on the lattice. Similarly, it is anticipated that the 1/3 and 5/9 plateaus also correspond to the MCS [16, 17, 24, 25]. The state of the 1/9 plateau remains theoretically contentious, predicted to be as either a $Z_{3}$ spin liquid [17] or a VBC [18, 21] state, with detailed magnetic structure yet to be elucidated. Very recently and interestingly, the 1/9 and 1/3 plateaus have also been experimentally observed in the kagome compound $\rm YCu_{3}(OH)_{6+\it x}Br_{3-\it x}$ ($x\approx 0.5$[26, 27, 28]. Therefore, theoretical and experimental studies on magnetization plateaus in the kagome lattice are currently attracting significant attention. These theoretical studies have primarily focused on the ideal kagome lattice where all exchange interactions are equivalent. However, in many kagome antiferromagnetic compounds, distorted kagome lattices are formed, resulting in non-equivalent exchange interactions [29]. Nevertheless, theoretical investigations of distorted kagome lattices are scarce [30, 31]. Therefore, it can be expected that new quantum states, which do not exist in undistorted kagome lattices, may emerge.
@@
$\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$ is a recently synthesized compound where copper ions ($\rm Cu^{2+}$) with $S=1/2$ spins form a distorted kagome lattice [32, 33, 34, 35, 36, 37]. There are three types of nearest-neighbor exchange interactions as shown in Fig. 1. A previous study has theoretically revealed the existence of a ground state with ${\mathbf{Q}}=(1/3,1/3)$, coplanar long-range order in the classical spin system under zero magnetic field, which has also been experimentally observed by means of neutron scattering experiments. The candidate values for the exchange interactions have been proposed in two previous studies [35, 37]. In this paper, we refer to them as Type 1 [35] and Type 2 [37]. However, the quantum magnetic properties of the spin system under a magnetic field have not been sufficiently investigated, and thus remain unclear. Therefore, it is important to investigate the ground state in the quantum spin model for $\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$.
@@
In this study, we investigate the ground state of this model under a magnetic field by exact diagonalization using the Lanczos method. We reveal that clear 1/3, 5/9, and 7/9 plateaus are present for both the Type 1 and Type 2 parameter sets. These ground states are identified as the MCS based on their spin structures. Given the uncertainty in the exchange interactions of $\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$, we perform additional calculations by varying the exchange interactions within a certain range. Consequently, we find additional plateau phases at $M/M_{\rm{sat}}=1/3$, 5/9, and 7/9, each distinctly different from the MCS. We argue that clear magnetization plateaus can be observed through high-field magnetization measurements conducted on $\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}$.
@@ II Model
$J_{\rm r}/J_{\rm b}$ 0.869 1
$J_{\rm g}/J_{\rm b}$ 0.056 0.45
We perform exact diagonalization using the Lanczos method for clusters with $N=27$, 36, and 45 under periodic boundary conditions (PBCs) as shown in Fig. 1, where $N$ represents the number of sites. Our calculations for $N=45$ cannot be performed some subspaces with $M=1/2,3/2,...,45/2$. For example, the $M=1/2$ subspace has the largest dimension of $\binom{45}{22}$, making the computation time and memory insufficient. Therefore, calculations are conducted within the range of $M\geq 23/2$ ($M/M_{\rm{sat}}\geq 23/45$), where the dimensions of these subspaces are $\binom{45}{11}$ or less.
In the following, we describe this model under four different limits. When $J_{\rm r}=J_{\rm g}=J_{\rm b}$, it corresponds to the undistorted kagome lattice. When $J_{\rm g}=J_{\rm b}=0$ and $J_{\rm r}>0$, the model consists of isolated hexamers connected by $J_{\rm r}$ bonds and monomers. When $J_{\rm r}=J_{\rm b}=0$ and $J_{\rm g}>0$, the model consists of isolated trimers connected by $J_{\rm g}$ bonds. When $J_{\rm g}=0$, $J_{\rm r}>0$ and $J_{\rm b}>0$ the model exhibits frustration due to the presence of 9-site triangles, in which $J_{\rm r}$ bonds connect three trimers formed by $J_{\rm b}$ bonds. This means that both type 1 and type 2 models exhibit frustration.
Figure 2: Magnetization curves of the kagome lattice with (a) the Type 1 and (b) Type 2 parameters. The blue dashed, red solid, and green solid lines represent the calculation results for $N=27$, 36, and 45, respectively. The black solid lines show the calculation results for the classical spin system at $N=5184$. Fractional values indicate the values of $M/M_{\rm sat}$. The horizontal dashed lines are reference lines added for clarity.
@@ III.1 Magnetization curve of $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$
Figure 2 depicts the magnetization curves at zero temperature of the Type 1 and Type 2 for ${\rm Y_{3}Cu_{9}(OH)_{19}Cl_{8}}$. Here, we set $g=2$. Both the Type 1 and Type 2 exhibit clear plateaus at $M/M_{\rm{sat}}=$ 1/3, 5/9, and 7/9. In the Type 1, the widths of these plateaus show almost no dependence on the number of sites $N$, indicating negligible finite-size effects. Thus, these plateaus persist even in the thermodynamic limit. In the Type 2, while the effects of size are negligible for the 5/9 and 7/9 plateaus, the width of the 1/3 plateau varies significantly at $N=27$ and $N=36$. However, the plateau width is broader for $N=36$ than for $N=27$, indicating its significant persistence even in the thermodynamic limit. The magnetic structures of all three magnetization plateaus are confirmed to be the MCS in Sec. III.2. Figure 2 shows that all of these magnetization plateaus can be observed through magnetization measurements on $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$. Notably, the 1/3 plateau is observable at approximately 150 T, and the possibility of observing all plateaus increases at a stronger field of approximately 300 T, achievable with state-of-the-art magnetic field generators [38, 39, 40].
In Fig. 2(a), the width of the 1/6 plateau at $N=36$ is comparable to that of the 1/3 plateau. Hence, the presence of the 1/6 plateau is also suggested. Since the $N=27$ system can only have values of $M/M_{\rm{sat}}=n/27$ ($n=1,3,5,...,27$), the 1/6 plateau cannot exist. However, the plateau at $M/M_{\rm{sat}}=5/27$, which is closest to 1/6, is not wide. Therefore, there are insufficient results to conclude the existence of the plateau. The 1/6 plateau requires the number of sites to be a multiple of 12, and since the model has 9 sites per unit cell, $N$ must be a multiple of 36. Therefore, it is currently impossible to evaluate it through diagonalization.
Both the Type 1 and Type 2 show a magnetic jump immediately after $M/M_{\rm{sat}}=1/6$ at $N=36$. This jump occurs from $M/M_{\rm{sat}}$=3/18(1/6) to $M/M_{\rm{sat}}$=5/18, skipping $M/M_{\rm{sat}}$=4/18, suggesting the possibility of a first-order field-induced quantum phase transition. However, at $N=27$, no jumps, such as from $M/M_{\rm{sat}}$=5/27 to $M/M_{\rm{sat}}$=9/27 skipping $M/M_{\rm{sat}}$=7/27, are observed. Thus, the jump observed at $N=36$ might be special due to finite-size effects. Further calculations with larger clusters are also necessary.
@@
Figure 4: Magnetization curves for the kagome lattice with the exchange interactions $J_{\rm r}$, $J_{\rm g}$, and $J_{\rm b}$, where $J_{\rm r}$ and $J_{\rm b}$ are fixed at 1, and $J_{\rm g}$ varies from 0 to 2. The blue dashed, red solid, and green solid lines represent the calculation results for $N=27$, 36, and 45, respectively. Fractional values indicate the values of $M/M_{\rm sat}$. The horizontal dashed lines are reference lines added for clarity.
@@ III.2 Spin structures of the magnetization plateaus
respectively, while the thickness of the lines corresponds to their magnitude. The diameter of the magenta lattice points corresponds to the value of $\Braket{S_{i}^{z}}$ at that site. The spin correlations between bonds connected by the exchange interaction $J_{\rm r}$ are strongest, forming periodically arranged hexagons. The remaining spins align parallel to the magnetic field. This structure corresponds to the MCS. Therefore, the magnetic structures of these plateaus are the same as the anticipated magnetic structures of the plateaus arising in the uniform kagome lattice [16, 17, 23, 24, 25]. The emergence of the 1/3, 5/9, and 7/9 plateaus can be understood by considering a model with only the $J_{\rm r}$ hexagon. In this model, magnetization plateaus occur at $M=0,1,2$. These plateaus, combined with the remaining spins aligned parallel to the magnetic field, give rise to the $M/M_{\rm sat}$=1/3, 5/9, and 7/9 plateaus on the kagome lattice. In the hexagon model, the exact values of the correlation $\Braket{\mathbf{S}_{i}\cdot\mathbf{S}_{j}}_{C}$ are -0.467 at $M=0$, -0.381 at $M=1$, and -0.194 at $M=2$. In the Type 2 model, the values of the correlation are -0.263 at $M/M_{\rm{sat}}=1/3$, -0.222 at $M/M_{\rm{sat}}=5/9$, and -0.103 at $M/M_{\rm{sat}}=7/9$, as shown in Fig. 3. These values in the Type 2 model are about 60% of those in the hexagon model. The presence of $J_{b}$ and $J_{g}$ weakens their correlations, but the correlation within the $J_{r}$ bonds remains significantly larger compared to those within the $J_{b}$ and $J_{g}$ bonds.
@@ III.3 Calculations with other parameters
Two sets of exchange interaction parameters, the Type 1 and Type 2, listed in Table 1, have been proposed for $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$. While these parameter sets are not significantly different, they do not completely match. Analyzing the model with wider range of parameters help us to get a deeper understanding of its magnetism. Moreover, in the future, when new compounds are synthesized and their magnetic measurements are conducted, analyzing it based on alternative parameters would provide significant assistance. This is important from a theoretical perspective as well, given the possibility of discovering new magnetic phases. In the Type 1, $J_{\rm r}\approx J_{\rm b}\neq J_{\rm g}$ while in the Type 2, $J_{\rm r}=J_{\rm b}\neq J_{\rm g}$ as shown in Table 1. Therefore, in this section, we investigate the model with the restriction $J_{\rm r}=J_{\rm b}\neq J_{\rm g}$. We consider the normalized Hamiltonian (1) with $J_{\rm r}=J_{\rm b}=1$ and the magnetic field term $h=g\mu_{\rm B}H$.
@@
Figure 4 shows the magnetization curves for $0\leq J_{\rm g}\leq 2$. Interestingly, even at $J_{\rm g}=0$, the 1/3, 5/9, and 7/9 plateaus are observed, as evident from Fig. 4(a). Furthermore, for $J_{\rm g}<1$, these plateaus exhibit a broader width compared to that for $J_{\rm g}=1$. This is because, in the region where $J_{\rm g}$ is small, $J_{\rm r}$ is relatively the largest, stabilizing the MCS, which is a state with strong correlations within the hexagons formed by $J_{\rm r}$ bonds. Considering that the clear 5/9 and 7/9 magnetization plateaus have not yet been observed experimentally in the kagome lattice, the theoretical prediction of increased plateau widths provides sufficient motivation to conduct further experiments. For $J_{\rm g}>1$, the 5/9 and 7/9 plateau widths narrow as $J_{\rm g}$ increases. At $J_{\rm g}=1.5$, the plateau widths are significantly narrowed, as shown in Fig. 4(g), suggesting that these plateaus may not exist in the thermodynamic limit. As $J_{\rm g}$ becomes larger, $J_{\rm r}$ becomes relatively smaller. Consequently, the MCS becomes unstable. Therefore, in the region where $J_{\rm g}$ is large, the 5/9 and 7/9 plateaus no longer exist. In contrast, the width of the 1/3 plateau widens as $J_{\rm g}$ increases. At $J_{\rm g}=2$, as depicted in Fig. 4(h), it is evident that the width of the 1/3 plateau significantly widens. This is a natural consequence of the fact that when $J_{\rm g}$ becomes very large, it becomes equivalent to an isolated three-site model, resulting in the broadening of the 1/3 plateau. Since in the region where $J_{\rm g}$ is large, the MCS becomes unstable, it is expected that a phase transition exists around $J_{\rm g}=1$. This will be discussed in detail later.
@@
To confirm the presence or absence of phase transitions with respect to $J_{\rm g}$ at $M/M_{\rm sat}=1/3$, 5/9, and 7/9, we calculate the second derivative of the energy per site $e=E/N$ with respect to $J_{\rm g}$. Figure 5 shows the calculation results of $-\frac{d^{2}e}{dJ_{\rm g}^{2}}$ at $M/M_{\rm sat}=1/3$, 5/9, and 7/9. It is evident that all three results exhibit sharp peaks at $J_{\rm g}=1$. For $M/M_{\rm sat}=1/3$ and 5/9, based on the results in Figs. 5(a) and 5(b), it is not possible to precisely determine the dimension of the phase transition, but it is expected to be either second-order or first-order. Further calculations with larger system sizes are required to clarify the dimension of the phase transition. For $M/M_{\rm sat}=7/9$, as shown in Fig.5(c), it is apparent that $-\frac{d^{2}e}{dJ_{\rm g}^{2}}$ diverges at $J_{\rm g}=1$ sharply, indicating a first-order phase transition at $J_{\rm g}$.
@@
Figure 6: Results of finite-size scaling for the 5/9 plateau at $J_{\rm r}=J_{\rm b}=1$ and $J_{\rm g}=1.2$. The lower black line and upper black line represent the results of the linear fits for $h_{\rm l}$ and $h_{\rm u}$, respectively, which are defined as the lower and upper magnetic fields of the magnetization plateau. The red and blue points represent $h_{\rm l}$ and $h_{\rm u}$ in the thermodynamic limit, respectively, and the bars represent the error from the linear fitting.
@@
In the following we discuss the existence of plateaus in the thermodynamic limit. The plateau width for $M/M_{\rm sat}=1/3$ remains broad regardless of $J_{\rm g}$, suggesting the existence of plateaus in the thermodynamic limit. In contrast, for $M/M_{\rm sat}=5/9$ and $7/9$, particularly at $J_{\rm g}>1$, the narrow plateau widths, as depicted in Fig. 4, make it difficult to discern the presence of plateaus in the thermodynamic limit. Therefore, it is necessary to examine finite-size effects on the plateau width. In such analyses, first, the existence of gapless excitations is assumed and an extrapolation is performed using the size dependence of $1/N$ or $1/N^{1/2}$ [41, 42], where the former is used in this study. Let us define the lower magnetic field and upper magnetic field of the magnetization plateau are defined as $h_{\rm l}$ and $h_{\rm u}$, respectively. If $h_{\rm u}-h_{\rm l}$ remains positive in the thermodynamic limit $N\rightarrow\infty$, it is interpreted as an evidence of the presence of plateaus [43, 44, 45]. Figure 6 shows the result of finite-size scaling for the 5/9 plateau at $J_{\rm g}=1.2$. The black solid lines at the lower and upper sides correspond to $h_{\rm l}$ and $h_{\rm u}$, respectively. The red and blue points represent $h_{\rm l}$ and $h_{\rm u}$ in the thermodynamic limit, respectively, and error bars from the linear fitting are also shown. Here, the errors for $h_{\rm l}$ and $h_{\rm u}$ are denoted as $\delta h_{\rm l}$ and $\delta h_{\rm u}$, respectively. The estimates of finite-size effects from linear fitting are not completely accurate when the system size is not large. Therefore, in this paper, we cautiously estimate that the plateau exists if the condition
@@
is satisfied. Table 2 and Table 3 present $W_{\rm min}$ for $M/M_{\rm sat}=5/9$ and 7/9, respectively, at various $J_{\rm g}$. For both $M/M_{\rm sat}=5/9$ and 7/9, $W_{\rm min}$ remains positive up to $J_{\rm g}=1.2$, indicating the presence of the plateaus for $J_{\rm g}\leq 1.2$.